OpenStudy (anonymous):

Diff Eq problem: \[\frac{ dy }{ dt } + t^2y = 1\] So, first I find my integrating factor. \[e ^{\int\limits_{}^{}t^2dt} = e ^{2t+c}\] Next, I multiply both sides by the integrating factor. \[(\frac{ dy }{ dt } + t^2y)e ^{2t+c} = e ^{2t+c} \] Then I want to integrate both sides so I should have : \[\int\limits_{}^{}d(e ^{2t+c} )y = \int\limits_{}^{} e ^{2t+c} dt\]

OpenStudy (anonymous):

Is this looking good so far @phi ?

OpenStudy (anonymous):

so now I've got \[e ^{2t+c}y = \frac{ 1 }{ 2 }e ^{2t+c} \] hmmmmm, at this point its not looking right, and not headed toward the answer in the book.

OpenStudy (anonymous):

I should have \[y = e ^{-1/3t^3} \int\limits_{}^{}e ^{-1/3t^3}dt+c\]

OpenStudy (inkyvoyd):

line 2 is incorrect

OpenStudy (inkyvoyd):

\(e ^{\int\limits_{}^{}t^2dt} = e ^{t^3/3+c}\)

OpenStudy (austinl):

I don't believe you need the "+C" in the exponential.

OpenStudy (inkyvoyd):

Austin is correct probably, seeing as you can multiply both sides by whatever expression you want as long as you don't multiply by zero and divide after

OpenStudy (austinl):

\(y\prime + t^2y=1\) \(\mu(t)=e^{\int p(t)dt}\) \(p(t)=t^2\) \(\int t^2dt=\dfrac{t^3}{3}\) \(\mu(t)y\prime+t^2\mu(t)y=\mu(t)\)

OpenStudy (inkyvoyd):

staph! I wanted the medal

OpenStudy (austinl):

Then I do believe that you simply integrate both sides and solve for y if feasible?

OpenStudy (anonymous):

Thanks a ton guys!

OpenStudy (austinl):

No problemo!